OpenStudy (anonymous):

Which ratio is equivalent to the ratio 2 to 10? 2 : 5 1 to 10 one over three 3 : 15

OpenStudy (anonymous):

2 to 5

OpenStudy (anonymous):

2 : 5

OpenStudy (e.mccormick):

\(2:10 = 1:5\) They work as fractions: \(\dfrac{2}{10}=\dfrac{1}{5}\) There is another one it is the same as that happens to be an answer, but it is NOT 2:5. \(\dfrac{2}{5}=\dfrac{4}{10}\ne\dfrac{2}{10}\)